OpenStudy (anonymous):

If 3^5=x, which of the following expressions is equal to 3^11

OpenStudy (anonymous):

243x 3x^2 9x^4 27x^3 x^6 @Directrix

OpenStudy (zzr0ck3r):

3^5=x 3^6*3^5=x*3^6 3^11=3^6 * x 3^11 = 729x

OpenStudy (zzr0ck3r):

hmm

OpenStudy (anonymous):

plug in 3^5 to ever x in your answer choices till you get 3^11

OpenStudy (zzr0ck3r):

3^5 = x square both sides (3^5)^2 = x^2 3^10=x^2 multiply by 3 3*3^10 = 3x^2 3^11 = 3x^2

Directrix (directrix):

729x = 3x^2 @zzr0ck3r I guess this is a problem in regrouping.

OpenStudy (anonymous):

@Directrix @timo86m @zzr0ck3r Thank you all :-) I have to go now.

Directrix (directrix):

Okay. Let's practice Geometry SAT tomorrow. @sakigirl

Directrix (directrix):

We should have a regular "Practice your SAT group" every night and do a few problems together. @Everybody

OpenStudy (zzr0ck3r):

im there

OpenStudy (anonymous):

@Directrix Agreed